# Residual subspace evolution strategies for nonlinear inverse problems

### v1

@@ 4.4 Nonlinear deconvolution
We reconstruct a length-$128$ signal $\mathbf{x}_{\star}$. We blur it densely, pass it through a saturating nonlinearity, and add noise. A second variant of this benchmark uses perturbed residual weighting to test robustness. The forward model reads
@@
It uses blur matrix $\mathbf{A}\in\mathbb{R}^{n\times n}$ and observations
@@
We resample the jitter $\nu$ on every call, introducing a small Monte Carlo perturbation atop the saturating nonlinearity. For each blur we draw $\mathbf{A}_{ij}\sim\mathcal{N}(0,1/n)$, average it with its transpose, add $3I_{n}$, and scale by $1/5$ to obtain a well-conditioned low-pass filter. We round outputs to $10^{-3}$ before adding noise and set
@@
With intact residuals (Fig. 4(a)), RSES achieves near-best state error and near-lowest residual, and finishes far faster than every baseline. EKI posts the smallest residual but at much higher computational cost. NEWUOA trails in both accuracy and runtime. xNES remains inaccurate, while GN runs quickly but stays far from the target.
@@
(b) Perturbed residual weighting with correlated components.
@@
(b) Perturbed residual weighting with correlated components.
@@
5 Conclusions and future work
Residual subspace evolution strategies address nonlinear inverse problems using residual-only information. Across deterministic, noisy, and non-differentiable benchmarks, RSES matched or exceeded every baseline. It solved the linear system to machine precision alongside Gauss-Newton, achieved the best residual and parameter accuracy in Brownian calibration while maintaining reasonable cost, attained the lowest losses on noisy MLP regression while running faster than Adam and far faster than NEWUOA, and ranked at or near the best errors on both deconvolution variants while running substantially faster than EKI.
NEWUOA and related model-based solvers excel on low-dimensional tasks but become costly because they require at least $O(n)$ evaluations per iteration. RSES gains advantage as dimension increases because it operates with $k\ll n$ probes. Its maintained accuracy under residual weight perturbations underscores robustness to model misspecification.
RSES depends on random probes that must explore useful directions. When $k\ll n$, poorly aligned samples can slow progress even with a fixed probe scale. Current parameter selection remains heuristic: we do not yet match ridge or probe-scale settings to local curvature or noise levels. Theoretical convergence guarantees under Monte Carlo perturbations remain open.
Future work will address these limitations through adaptive ridge tuning, preconditioned or structured probe sampling, and parallel residual evaluations. We plan to test RSES on larger imaging and simulation workloads without adjoints, probe objectives with many local minima to assess global robustness, and combine RSES with coarse surrogate models or reduced-order bases when random probes alone cannot efficiently cover the full state space.

### v2

@@ 3.4 Nonlinear deconvolution
The final benchmark reconstructs a length-$128$ signal $\mathbf{x}_{\star}$ from blurred, saturated, and noisy observations. The forward model applies dense blur, passes the result through a saturating nonlinearity, and adds noise:
@@
A blur matrix $\mathbf{A}\in\mathbb{R}^{n\times n}$ generates the observations
@@
The jitter $\nu$ resamples on every call, introducing a small Monte Carlo perturbation atop the saturating nonlinearity. For each blur the entries of $\mathbf{A}$ are drawn from $\mathcal{N}(0,1/n)$, averaged with the transpose, shifted by $3I_{n}$, and scaled by $1/5$ to obtain a well-conditioned low-pass filter. Outputs are rounded to $10^{-3}$ before adding noise, with
@@
Intact residuals.
With the intact weighting (fig. 4(a)), RSES achieves near-best state error and near-lowest residual while finishing far faster than every baseline. EKI posts the smallest residual but at much higher computational cost. NEWUOA trails in both accuracy and runtime. xNES remains inaccurate, while GN runs quickly but stays far from the target.
Perturbed residuals.
@@
(b) Perturbed residual weighting.
@@
(b) Perturbed residual weighting.
@@
4 Conclusion
Residual subspace evolution strategies address nonlinear inverse problems using residual-only information. Across deterministic, noisy, and non-differentiable benchmarks, RSES matched or exceeded every baseline: it solved the linear system to machine precision alongside Gauss-Newton, achieved the best residual and parameter accuracy in Brownian calibration while maintaining reasonable cost, attained the lowest losses on noisy MLP regression while running faster than Adam and far faster than NEWUOA, and ranked at or near the best errors on both deconvolution variants while running substantially faster than EKI.
NEWUOA and related model-based solvers excel on low-dimensional tasks but become costly because they require at least $O(n)$ evaluations per iteration. RSES gains an advantage as dimension increases because it operates with $k\ll n$ probes. Its maintained accuracy under residual weight perturbations underscores robustness to model misspecification.
The method depends on random probes that must explore useful directions. When $k\ll n$, poorly aligned samples can slow progress even with a fixed probe scale. Current parameter selection remains heuristic: the ridge and probe-scale settings do not yet adapt to local curvature or noise levels. Theoretical convergence guarantees under Monte Carlo perturbations remain open.
Future work will address these limitations through adaptive ridge tuning, preconditioned or structured probe sampling, and parallel residual evaluations. Plans include testing RSES on larger imaging and simulation workloads without adjoints, probing objectives with many local minima to assess global robustness, and combining RSES with coarse surrogate models or reduced-order bases when random probes alone cannot efficiently cover the full state space.
@@ Appendix A Idealised descent on the probe span
Appendix B Ridge parameter selection
This appendix derives the residual-dependent ridge schedule $\lambda_{t}=\max(\beta\lVert\mathbf{r}_{t}\rVert_{2}^{2},\,\epsilon)$ used in the main algorithm.
Uniform bound on the ridge solution.
The ridge problem (7) has the closed-form minimiser
$\mathbf{w}^{*}=-(\mathbf{B}^{\top}\mathbf{B}+\lambda I)^{-1}\mathbf{B}^{\top}\mathbf{r}_{t}.$ (40)
Let $\mathbf{B}=\mathbf{U}\boldsymbol{\Sigma}\mathbf{V}^{\top}$ be a singular value decomposition. Then
$(\mathbf{B}^{\top}\mathbf{B}+\lambda I)^{-1}\mathbf{B}^{\top}=\mathbf{V}(\boldsymbol{\Sigma}^{2}+\lambda I)^{-1}\boldsymbol{\Sigma}\mathbf{U}^{\top},$ (41)
and the operator norm satisfies
$\lVert(\mathbf{B}^{\top}\mathbf{B}+\lambda I)^{-1}\mathbf{B}^{\top}\rVert_{2}=\max_{\sigma\geq 0}\frac{\sigma}{\sigma^{2}+\lambda}.$ (42)
The function $f(\sigma)=\sigma/(\sigma^{2}+\lambda)$ attains its maximum at $\sigma=\sqrt{\lambda}$, where $f_{\max}=1/(2\sqrt{\lambda})$. Hence
$\lVert\mathbf{w}^{*}\rVert_{2}\leq\frac{1}{2\sqrt{\lambda}}\lVert\mathbf{r}_{t}\rVert_{2}$ (43)
for any matrix $\mathbf{B}$, and the bound is tight.
Residual-dependent schedule.
Choosing $\lambda\geq\beta\lVert\mathbf{r}_{t}\rVert_{2}^{2}$ with $\beta>0$ yields
$\lVert\mathbf{w}^{*}\rVert_{2}\leq\frac{1}{2\sqrt{\lambda}}\lVert\mathbf{r}_{t}\rVert_{2}\leq\frac{1}{2\sqrt{\beta}},$ (44)
so the coefficient norm remains uniformly bounded regardless of the residual magnitude or the structure of $\mathbf{B}$. To prevent vanishing regularisation when the residual becomes small, the algorithm sets
$\lambda=\max\{\beta\lVert\mathbf{r}_{t}\rVert_{2}^{2},\,\epsilon\},$ (45)
which ensures
$\lVert\mathbf{w}^{*}\rVert_{2}\leq\min\!\left\{\frac{1}{2\sqrt{\beta}},\,\frac{\lVert\mathbf{r}_{t}\rVert_{2}}{2\sqrt{\epsilon}}\right\}.$ (46)
The first term caps the coefficient norm when the residual is large; the second keeps the step size proportional to the residual once it falls below $\sqrt{\epsilon/\beta}$. This adaptive schedule tightens the ridge as the iterate improves while maintaining numerical stability near convergence.